oth), wo - ) Problem 1. In this problem, we use to the to--) to approximate f'(o), i.e. f'(30) (Hint: Use the Taylor expansions for both f(+ h) and f(xo-h)). (A). Show, using the Taylor expansions for both f(10+h) and f(10-h), that the leading term of the error is "( 70) |h2.
Answer #1

Using Taylors Series Expansion. We know that ++++h) = 4(4) + h 460) + Sum) + n° {cx.) + bil ps3 *** 4670-) = 467.) - h f(f(xoth) - f(xo-h) - (x) + f (x0) + ... 2h 5! So, Error to approximate f(xo) by finite difference is of orden i f(xoth)-f(xo-

I hope it helps. Please feel free to revert back with further queries.